OpenStudy (anonymous):

(EASY 5TH GRADE MATH! MEDAL AND FAN!) Marta place 4 marbles in each of 3 bags. Then she placed 3 marbles in each of 4 bags. She noticed that both groups has 12 marbles. Which property of multiplication does this represent? A) Associative Property of Multiplication B) Commutative Property of Multiplication C) Identity Property of Multiplication D) Zero Property of Multiplication

OpenStudy (anonymous):

@ganeshie8 @dan815 @phi @whpalmer4

OpenStudy (anonymous):

I think it is A.

OpenStudy (anonymous):

Okay does everyone agree?

OpenStudy (phi):

It is a strange way to show 3*4 = 4*3

OpenStudy (nikkibxox):

C, I think because it stays 12, but who am I. ^-^

OpenStudy (anonymous):

PERMUTATION

OpenStudy (anonymous):

So what is it then?

OpenStudy (danjs):

look up the meaning of each of them

OpenStudy (anonymous):

Okay does everyone agree C? @phi

OpenStudy (anonymous):

Okay thank you everyone!

OpenStudy (danjs):

no

OpenStudy (anonymous):

No?

OpenStudy (phi):

commutative property

OpenStudy (anonymous):

Why is that?

OpenStudy (danjs):

3 * 4 = 12 4 * 3 = 12 3*4 = 4*3

OpenStudy (danjs):

http://www.mathsisfun.com/associative-commutative-distributive.html look at the meaning of it

OpenStudy (danjs):

identity has to do with multiplying by 1 zero has to do with multiplyhing by zero look at the forms of communative and associative , it is one of them ,